# Application of Machine Learning in Agriculture: Recent Trends and Future Research Avenues

### v1

@@ 5.4 Machine Learning Application in Disease Detection
Phenotypying 2019 DL, CNN CNN based modeling is used in computer vision for automated plant phenotyping, enabling high-throughput analysis of plant characteristics. [123]
@@ 6 Statistical and Bibliometric Analysis
In this section, we present the results of the statistical analysis conducted to comprehend the extent of research activity in the field of ML for agriculture. Statistical data has been gathered, revealing a substantial growth trend in recent years, highlighting it as one of the most dynamic and vibrant research domains.
Figure 7: The amount of research carried out in machine learning for agriculture. Data source: Scopus on May 19, 2023.
Fig. 7 showcases the volume of research conducted in the intersection of machine learning (ML) and agriculture. The search criteria encompassed the terms ‘machine learning’ and ‘agriculture’ in the title, abstract, or keywords of the publications. Each bar signifies a distinct year, with its height denoting the count of research articles during that year. A perusal of the bars highlights the evolving trajectory of research in this domain. A marked surge is evident between 2018 and 2020. An exponential trendline is plotted, boasting an $R^{2}$ value of $0.989$. This coefficient of determination indicates how well the trendline fits the actual data. An $R^{2}$ value nearing 1, as seen here, suggests that the trendline provides an exceptional fit, underscoring a robust correlation between years and publication count. The near-perfect fit of the exponential trendline, accompanied by the high $R^{2}$ value, reiterates the burgeoning research momentum in ML applied to agriculture. The bar chart succinctly captures this escalating research enthusiasm over the years, highlighting the field’s rising prominence.
Figure 8: Research contribution by countries in machine learning for agriculture. Data source: Scopus on May 19, 2023.
Fig. 8 presents the research contribution of different countries in ML for agriculture. The pie diagram visually represents the distribution of research publications, with each slice corresponding to a specific country or the ‘Others’ category. The size of each slice represents the proportionate contribution of that country to the total number of publications. India has the largest slice, indicating the highest number of research publications among the listed countries, followed closely by the United States and China. This pie diagram demonstrates a clear overview of the relative contributions of different countries in the field of ML for agriculture.
Table 7: Publication distribution by document Type. Data source: Scopus on May 16, 2023.
@@
In addition, Table 7 provides the distribution of publications by document type in the field of ML for agriculture. Moreover, the table provides the percentage distribution of each document type, which offers a relative perspective on their contribution to the total number of publications. For instance, articles accounted for $49.63\%$ of the total publications, while conference papers constituted $37\%$. The remaining document types had smaller percentages, ranging from $0.20\%$ for editorials to $4.45\%$ for book chapters. In summary, the table offers a concise overview of the publication distribution in ML for agriculture, shedding light on the prevalence of various document types and their respective proportions within the field.
@@
Figure 9: Number of publications by different types of publishers. Data source: Scopus on May 19, 2023.
(a) Publication in IEEE
(b) Publication in Elsevier
(c) Publication in Springer
(d) Publication in MDPI
(e) Publication in miscellineous
(f) Publication in conferences
Figure 10: Distribution of articles on the intersection of ML and agriculture across various journals.
The contribution of publishers is depicted in Fig. 9. The majority of papers are published by Elsevier, followed by Mdpi, Springer Nature, Wiley, IEEE, and others. More detailed statistics for publication house are provided in Fig. 10. This statistical analysis provides valuable insights into the volume of research conducted and the distribution of publications in ML for agriculture. The substantial growth trend observed in recent years reinforces the dynamic and vibrant nature of this research domain.
@@
Figure 11: The bibliometric analysis of keywords in ML application in agriculture. Data source: Web of Science on May 19, 2023.
The bibliometric analysis of keywords (Fig. 11) in ML application in agriculture provides valuable insights into the key areas of focus within this field. In this paper, the open-source application VoSviewer [124] has been used for bibliometric analysis. By examining the frequency and trends of specific keywords, researchers can identify the most prominent topics and research directions. The analysis reveals that keywords such as ‘precision farming’, ‘crop yield prediction’, ‘plant disease detection’, and ‘soil analysis’ are among the most frequently mentioned in research publications. This indicates a strong emphasis on utilizing ML techniques for optimizing farming practices, improving crop productivity, and addressing agricultural challenges. By understanding the prevalent keywords, researchers and practitioners can gain a better understanding of the areas where ML is making significant contributions and guide future research efforts.
@@
Figure 12: The bibliometric analysis of countries in ML application in agriculture. Data source: Web of Science on May 19, 2023.
The bibliometric analysis of countries (Fig. 12) in ML application in agriculture offers a comprehensive view of the research contributions from different nations. By assessing the number of research publications, researchers can identify the countries that are leading in this field. The analysis reveals that India, the United States, and China are among the top contributors in terms of research publications. These countries demonstrate a strong commitment to leveraging ML techniques for agricultural applications and have made substantial contributions to the advancement of the field. Additionally, the analysis highlights the contributions of other countries collectively categorized as ”Others,” indicating a global interest in ML application in agriculture. By recognizing the leading countries, researchers can foster collaborations and knowledge-sharing initiatives to further accelerate the progress in ML for agricultural innovation.
@@
Figure 13: The bibliometric analysis of organisations in ML application in agriculture. Data source: Web of Science on May 19, 2023.
The bibliometric analysis of organisations in ML application in agriculture is shown in Fig. 13. The advancement of scientific research relies on the collaborative efforts of numerous institutions worldwide. Aarhus University, Addis Ababa Science and Technology University, and Agricultural Research Center are just a few examples of the organizations contributing to the knowledge base. These institutions have produced a significant number of research documents and garnered citations, demonstrating their impact on their respective fields. With a strong link strength, they have fostered connections and collaborations within the scientific community, further enhancing the progress of scientific inquiry.
7 Case Studies and Examples

### v2

@@ 3.3 Machine Learning Applications in Post-Harvesting
CNN-based modeling is used in computer vision for automated plant phenotyping, enabling high-throughput analysis of plant characteristics. [207]
3.4 Machine Learning Application in Disease Detection
Researchers have extensively used ML and image processing techniques to identify diseases in various types of plants during pre-harvest, harvest, and post-harvesting. Hence, a separate discussion is required to address this focus.
In [153] an automated approach is proposed for the detection of leaf disease in agriculture for tomato plants. The model aims to detect diseases in early stages, enhancing crop production. SVM, Random Forest, and Logistic Regression algorithms are compared, and SVM outperforms the others. The study lacks specific details on the size of the data set and the evaluation metrics. Similarly, a system for the detection of apple disease based on ML and computer vision is proposed in [154]. It achieves $98.63\%$ accuracy by segmenting diseased regions and utilizing features such as DWT and color histograms.
For accurate and efficient multi-class plant disease detection, a deep learning object detection model is proposed in [155], surpassing the performance of existing models with a mean average precision of $91.2\%$ and an F1-score of $95.9\%$. Another study [156] introduces a deep convolutional neural network (DCNN) for early diagnosis of apple tree leaf diseases (ATLDs). The proposed DCNN achieves $98.82\%$ accuracy, showing fast convergence, fewer parameters, and high robustness. In [157], an accurate identification approach for common apple leaf diseases is proposed using a deep convolutional neural network. It achieves a $97.62\%$ accuracy with reduced parameters. A lightweight convolutional neural network, RegNet, is introduced in [160] for rapid and accurate identification of apple leaf diseases, surpassing other pre-trained models with a $99.23\%$ overall accuracy. The research in [159] presents a CNN model with $19$ layers for accurate classification of Marsonina Coronaria and Apple Scab diseases in apple leaves, achieving $99.2\%$ accuracy on a dataset of $50,000$ images. Another ML-based approach for the automated detection of apple tree diseases is proposed in [158], achieving a high accuracy of approximately $95\%$ for the classification of diseases. A deep convolutional neural network model is proposed in [161] for real-time detection of apple leaf diseases, achieving an overall accuracy of $99.31\%$ and surpassing other models. The research in [162] utilizes fine-tuned pre-trained CNN models for accurate apple crop disease detection, with DenseNet201 achieving $98.75\%$ accuracy.
In [163], an automated software solution is proposed using image processing techniques for the detection of apple leaf disease, based on color and texture characteristics. Finally, an algorithm for accurate segmentation of diseased parts in apple leaf images is proposed in [164], achieving an accuracy of $96.4\%$ with the $k$ nearest-neighbor classifier. A detailed look is given in Table 4.
Table 4: Disease Detection Applications in Agriculture (Pre-harvesting and Post-harvesting)
Sn Application Year Features Models Ref.
1 Pre-harvesting disease detection 2020 CNN is used to automatically identify crop diseases $\bullet$ Inception-ResNet-v2 [134]
2 Pre-harvesting disease detection and crop recommendation system 2021 Presents a location and season-based crop recommendation system. $\bullet$ CNN
$\bullet$ VGG-16
[137]
3 Pre-harvesting disease detection 2021 A hybrid system based on a Convolutional Autoencoder (CAE) network and CNN for automatic plant disease detection. $\bullet$ CNN [131]
4 Pre-harvesting disease detection 2021 Uses computer vision and AI for early detection of plant diseases, reducing continuous human monitoring challenges. $\bullet$ CNN [133]
5 Pre-harvesting disease detection 2022 A model using CNN and VGG to accurately classify disease-affected leaves. $\bullet$ CNN
$\bullet$ VGG
[132]
6 Pre-harvesting disease detection 2022 Employs a hybrid ML model and a web-based AI tool for plant disease identification. $\bullet$ CNN
$\bullet$ AlexNet
$\bullet$ LeNet
[135]
7 Post harvesting disease detection 2011 Automatically extracts quality features from external data using image processing. $\bullet$ CNN [73]
8 Post harvesting disease detection 2017 Incorporates a prediction system for warehouse managers to recommend optimal dispatch sequences and transportation conditions. $\bullet$ IoT
$\bullet$ ML
[74]
9 Post harvesting disease detection 2019 Uses image processing techniques to classify mangoes based on their condition across three phases. $\bullet$ ML
$\bullet$ DOF
$\bullet$ Edge Detection
$\bullet$ RGB
[70]
10 Post harvesting disease detection 2019 Detects calyx and stalk scars in both defected and healthy tomatoes. $\bullet$ RBF SVM [72]
11 Post harvesting disease detection 2023 Introduces a knowledge-based expert system utilizing a Bayesian network for diagnosing post-harvest diseases in apples. $\bullet$ Bayesian network [69]
3.5 Model used in Literature
To provide deeper insights into the utilization of ML models in literature, here is a detailed analysis of models applied across various domains from Table LABEL:tab:long and 3.
Neural networks (NN) and SVM are widely applied for soil classification [120] and weather prediction [79], offering significant accuracy and efficiency. However, their success depends heavily on parameter tuning and data quality. ML models integrated with IoT and wireless sensor networks have optimized irrigation planning, offering real-time monitoring and efficient water usage. Similarly, crop recommendation systems employ ML models like RF, SVM, and deep learning to provide farmers with practical decision-making tools, considering soil type and environmental factors [234]. However, issues such as sensor reliability and maintenance costs remain challenges.
CNNs and computer vision have been extensively used for detecting plant diseases and classifying seeds with high accuracy [132]. Advanced architectures like Inception-ResNetV2 and MobileNetV2 have further enhanced performance [134, 126]. In the domain of fruit detection and harvesting, models like YOLO and R-CNN have been pivotal in automating the tasks of detection in image data, reducing labor dependency, and improving efficiency [57, 167]. LSTM models have been mostly used for analyzing and predicting crop yield [65].
ML models have accelerated crop improvement by analyzing genetic data, predicting seed yield, and enabling precision breeding. Random Forest and Neural Networks, in particular, have proven effective in estimating phenotypic traits and improving crop yield [199]. Additionally, explainable AI is emerging in plant breeding to enhance model interpretability and decision-making [198].
The analysis highlights that CNNs dominate in image-based applications such as disease detection and fruit classification, while IoT-integrated models are prevalent in irrigation and post-harvest management. Traditional models like SVM and Random Forest are widely used due to their simplicity and efficiency, particularly for soil and weather-related applications. These findings emphasize the importance of selecting appropriate ML models for specific agricultural challenges while addressing their limitations to maximize their practical applicability and impact on sustainable farming practices.
Figure 2: Sankey diagram illustrating the integration of various ML models and technologies (e.g., IoT, YOLO) across different stages of agricultural processes, including pre-harvest, post-harvest, harvest handling, and plant breeding, contributing to improved food security and sustainable farming practices.
A concise visualizes the application of ML models across agricultural phases, based on insights from the literature (Table LABEL:tab:long and 3) is given in the sankey diagram Fig.2. From the diagram, it is clear that, in the pre-harvesting phase, ML models like CNN, SVM, and XGBoost are widely used for soil classification, seed testing, irrigation planning, and disease detection. These models support precision farming by analyzing soil properties, optimizing water use, and identifying crop diseases early. Harvesting relies heavily on computer vision models such as YOLO, R-CNN, and LSTM for fruit detection, picking, and yield prediction, ensuring automated and efficient crop assessment.
Post-harvesting integrates IoT, Bayesian Networks, and DL models for disease detection, food security, and handling, focusing on supply chain efficiency and quality control. Plant breeding involves phenotyping and climate-resilient breeding, leveraging Neural Networks, SVM, and DL to enhance crop adaptability and productivity.
3.5.1 Comparison of Machine Learning Techniques in Agriculture
ML models are widely used in agriculture for image processing, classification, prediction, and automation. Different ML techniques are suited to specific tasks. This section consolidates ML methods to avoid redundancy in subsections 3.1, 3.2, 3.3.
  • Convolutional Neural Networks (CNNs): CNNs are used for tasks where image data is being used such as crop disease detection, fruit classification, and yield estimation. CNNs provide high accuracy in feature extraction but require large datasets and high computational power [137] .
  • Support Vector Machines (SVMs): In ML classification tasks, including soil type identification and pest detection. SVMs work well with structured datasets but struggle with complex, high-dimensional data [129, 55].
  • K-Nearest Neighbors (K-NN): Used in soil moisture prediction and crop classification due to its ease of implementation. However, it becomes computationally expensive for large datasets [129].
  • Decision Trees & Random Forest: These ensemble learninf based algorithms re effective for yield prediction and irrigation optimization. These models handle missing data well and provide interpretable decision-making [56].
  • Extreme Learning Machines (ELM): ELMs are mostly applied in soil temperature estimation and regression-based problems due to their fast training speed [43] .
  • Deep Learning & Hybrid Models: These are advanced architectures that integrate CNNs, RNNs, and transformers for real-time agricultural monitoring and automation [25].
4 Information Fusion in Agriculture
To improve the accuracy of agricultural practices, researchers and practitioners integrate data from diverse sources—a process known as information fusion in agriculture [37]. Addressing RQ-2, this section examines common dataset sources and the types of datasets used in agricultural applications. It further discusses how ML-driven data fusion enhances agricultural outcomes through ensemble learning, hybrid models, data aggregation strategies, and transfer learning.
A database search in Scopus was performed on 30 December 2024 using the keywords “machine learning”, “data fusion”, and “agriculture” (searched within the title, abstract, or keywords). The resulting trends, depicted in Fig. 3, reveal a steady year-by-year increase in research on data fusion.
Figure 3: Trends in Multi-Source Data Fusion for Agricultural Machine Learning. Data source: Scopus, 30 December 2024.
4.1 Common Data Types in Agriculture
The common data types utilized in agriculture with ML, including RGB images, hyperspectral images, multispectral images, and text data, are summarized in Table 5. For instance, the study by Maimaitijiang et al. [42] evaluates the effectiveness of UAV-based multimodal data fusion—integrating RGB, multispectral, and thermal sensors—for estimating soybean (Glycine max) grain yield within a Deep Neural Network (DNN) framework. The researchers collected RGB, multispectral, and thermal images using a cost-effective multi-sensor UAV at a test site in Columbia, Missouri, USA. The study highlights three key findings:
  • Multimodal data fusion enhances yield prediction accuracy and improves adaptability to spatial variations.
  • DNN-based models outperform traditional methods, with the DNN-F2 model achieving the highest performance ($R^{2}=0.720$ and RMSE= 15.9%).
  • DNN models demonstrate greater resilience to saturation effects and exhibit more adaptive performance across different soybean genotypes.
Remote sensing data plays a critical role in precision agriculture, with key sources of data acquisition including satellite imagery, aerial data, drone (UAV) data, and LiDAR & radar data [37] . These data types are widely used for crop monitoring, disease detection, and yield prediction through multispectral, hyperspectral, and thermal imaging. Time-series data, primarily collected from IoT sensors, is crucial for monitoring environmental parameters such as temperature, humidity, and soil moisture [38]. Tabular data, consisting of structured information like soil composition, crop characteristics, and historical weather records, supports predictive analytics for yield estimation and market forecasting [39]. Text data, derived from farmer surveys and agricultural reports, helps in shaping more sustainable policies and providing farmers with accurate feedback, as seen in AI-based agricultural chatbot Kisan GPT [40]. Additionally, biological data, including genetic and phenotypic information, is essential for plant breeding, predicting plant traits, and estimating phenotypic characteristics for improvement of different crop varieties [198].
Table 5: Common Data Types and Data Modalities Used in Agriculture with Machine Learning
Data Category and Subcategory Data Modality Ref.
Remote Sensing Data Satellite Data Image Data (Multispectral, Hyperspectral, Thermal, Radar) [37]
Aerial Data Image Data (RGB, Thermal)
Drone (UAV) Data Image Data (RGB, Multispectral, Hyperspectral, Thermal)
LiDAR & Radar Data No Traditional Images (3D Point Clouds & Reflectance Data)
IoT Sensors with Spectral Imaging Image Data
Time-Series Data IoT Sensors Data Numerical Data [38]
Tabular Data Numerical & Categorical Data [39]
Text Data Text Data [40]
Biological Data (Genetic Data, Phenotypic Data) Image Data (RGB, Hyperspectral, Thermal), Tabular & Numerical Data [198]
4.2 Common Machine Learning Techniques for Data Fusion
  • Ensemble Learning for Multi-Source Data Integration: Ensemble learning methods, combines multiple models to improve prediction accuracy and robustness. Techniques such as Bagging, Boosting, and Stacking, help handle multi-source data by reducing individual model biases and improving the generalizability of predictions by ML models [21]. In the paper [20], the authors have used ensemble learning methods for yield prediction with (light detection and ranging) LiDAR and Multispectral imagery data from UAV (unmanned aerial vehicles). While in the paper [22], Extreme Gradient Boosting (XGBoost) has been used in monitoring agricultural soil organic carbon (SOC) with Sentinel-1 (S1) C-band dual polarimetric synthetic aperture radar (SAR) and Sentinel-2 (S2) multispectral data.
  • Bayesian Inference for data fusion: One of the common method for multi sensor data fusion is Bayesian theory that helps make better decisions on bridges. It is also used to image fusion of same scene into a single image that contains all the imporatnt features from each of the original images. The resulting images will be more suitable for machine leanring models for further decision making [33]. In the paper [23], the authors have proposed seven robust Bayesian optimized deep hybrid learning models for the automated classification of ten types of tomato leaves with testing of popular Plantvillage datasets.
  • Hybrid Models for Cross-Modal Learning: When multiple ML or deep learning (DL) architectures are combined to handle data from different sources, the resultant model is called a multimodal hybrid ML model. These methods help extract more features from multi-modal data, and the combined features are known as hybrid features. However, selecting the optimal fusion techniques for creating hybrid models remains critical for achieving optimal model performance [25]. The study [24] proposed a hybrid model named ITF-WPI for multimodal pest identification and classification in wolfberry. This model utilizes multimodal data (images and text), incorporating Transformer-based contextual extraction (CoT) and LSTM models for improved performance.
  • Data Aggregation Strategies for Large-Scale Agricultural Data: Data aggregation techniques merge data from multiple sources, such as sensor networks, satellite imagery, and IoT devices, into a unified format for ML models. Aggregating this type of data is crucial, as it becomes large-scale and complex, requiring structured processing for informed decision-making. These strategies include hierarchical aggregation, spatial aggregation and temporal aggregation , explain below briefly [27]. In this paper [28], the authors propose an efficient and scalable framework for privacy-preserving data aggregation in smart agriculture, utilizing data from smart devices and IoT networks. However the paper [26], explains a data aggregation scheme based on the ElGamal Cryptosystem ensuring, secure and private data processing of agricultural data taken with the use of IoT and 5th generation wireless network (5G).
  • Deep Learning-based Feature Fusion and Transfer Learning for Cross-Modal Data Adaptation: Deep learning-based feature fusion involves techniques that automatically extract and combine features from multiple sources, such as images, text, or sensors, using deep neural networks to improve predictive performance and enhance decision-making [34].
    In addition, transfer learning for cross-modal data adaptation allows ML models to leverage knowledge from one data modality (e.g., satellite imagery) and apply it to another (e.g., IoT sensor readings or time-series data). By fine-tuning pre-trained models for new tasks, this approach facilitates cross-domain adaptation and boosts performance in specific applications, such as precision agriculture, environmental monitoring, and smart city infrastructure [29]. For example, in the study [30], a general framework has been proposed to extract high-level latent feature representations using transfer learning methods.
4.3 Applications of Data Fusion and Hybrid Models in Various Agricultural Domains
This section highlights successful studies utilizing data fusion and hybrid ML models across various agricultural domains. While advanced fusion techniques are rapidly evolving in agriculture [31, 32], the literature still lacks extensive research in this area. This section discusses recent studies that incorporate data fusion and hybrid ML/DL models for applications such as crop monitoring, disease detection, yield prediction, crop identification and crop classification, and land monitoring in the Table 6.
Table 6: Some Studies on Data Fusion and Hybrid Models in Various Agricultural Domains
ML Technique Data / Sensor Used Application Areas Data Fusion Hybrid Models Ref.
$\bullet$ Partial Least Squares (PLS) Regression-Based Feature Fusion
$\bullet$ Ensemble Learning Model
PlantVillage Data Crop Disease Recognition $\times$ [41]
$\bullet$ Deep Neural Networks RGB, Multispectral, and Thermal Sensor Data Soybean Yield Prediction Using Low-Cost Sensors $\times$ [42]
$\bullet$ Extreme Learning Machine (ELM) Multispectral Data Grape Leaf Moisture Prediction [43]
$\bullet$ Transformer-Based GANet Model RGB, Thermal Infrared, and Depth Point Cloud Data Fusion Grassland Monitoring and Protection $\times$ [44]
$\bullet$ Laplacian Pyramid Transform (LPT)
$\bullet$ Fuzzy Logic
RGB and Thermal Images Fruit Detection [45]
$\bullet$ Backpropagation Neural Network
$\bullet$ Independent Recurrent Neural Network
Meteorological and Area Data Rice Crop Yield Prediction [46]
$\bullet$ Super-Resolution Convolutional Neural Network (ESRCNN) Landsat-8 and Sentinel-2 Data Sunflower Lodging Identification $\times$ [47]
@@ 5 Statistical and Bibliometric Analysis
Addressing RQ-3, this section outlines the findings of a statistical analysis conducted to evaluate research activity in ML for agriculture. The collected data indicate a substantial growth in research output in recent years, underscoring the dynamic and vibrant nature of this domain.
Fig. 4 illustrates the annual volume of publications at the intersection of ML and agriculture. The search criteria, which included the terms ‘machine learning’ and ‘agriculture’ in the title, abstract, or keywords, produced distinct yearly counts. A notable surge is observed between 2018 and 2020. An exponential trendline—with an $R^{2}$ value of 0.993—demonstrates an exceptional fit to the data, reflecting a robust correlation between publication count and time. This near-perfect correlation reinforces the rapidly escalating research momentum in ML applications for agriculture.
20132014201520162017201820192020202120222023202450000100000150000200000250000YearPublicationsPublicationsTrend Line $22444e^{0.212x}\times R^{2}=0.993$
Figure 4: The Amount of Research Carried Out in ML for Agriculture with Trend line. Data source: Scopus on December 30, 2024
Table 7 provides the distribution of publications by document type, along with their corresponding percentage shares. Articles comprise 49.63% of the total, while conference papers account for 37.00%. The remaining document types contribute smaller percentages, ranging from 0.20% for editorials to 4.45% for book chapters. This table succinctly captures the publication landscape within ML for agriculture.
Fig. 5 presents the research contributions by various countries. The pie chart clearly shows the proportional distribution of publications, with India leading, followed by the United States and China. This visual representation offers a concise overview of global research involvement in the field.
Table 7: Publication Distribution by Document Type. Data Source: Scopus on December 30, 2024
@@
Figure 5: Research Contribution by Countries in ML for Agriculture. Data Source: Scopus on December 30, 2024
Fig. 6 depicts the number of publications by different types of publishers, showing that Elsevier is the leading publisher, followed by MDPI, Springer Nature, Wiley, IEEE, and others.
@@
Figure 6: Number of Publications by Different Types of Publishers. Data Source: Scopus on December 30, 2024
Fig. 7 presents a bibliometric network of keywords in ML applications for agriculture, generated using VoSviewer [195]. Keywords such as ‘precision farming’, ‘crop yield prediction’, ‘plant disease detection’, and ‘soil analysis’ emerge as prominent, reflecting the central research themes in the field.
@@
Figure 7: The Bibliometric Analysis of Keywords in ML Application in Agriculture. Data Source: Web of Science on December 30, 2024
Fig. 8 offers a bibliometric mapping of countries contributing to ML research in agriculture. The analysis identifies India, the United States, and China as leading contributors, with additional nations grouped under ”Others,” underscoring the global reach of research in this area.
@@
Figure 8: The bibliometric analysis of countries in ML application in agriculture. Data source: Web of Science on December 30, 2024
Fig. 9 illustrates the bibliometric analysis of organizations in ML for agriculture. Institutions such as Aarhus University, Addis Ababa Science and Technology University, and Agricultural Research Center have emerged as key contributors, highlighting their impact and collaborative linkages within the research community.
@@
Figure 9: The Bibliometric Analysis of Organizations in ML Application in Agriculture. Data Source: Web of Science on December 30, 2024
Collectively, these statistical and bibliometric analyses provide a comprehensive overview of the research landscape in ML for agriculture. The rapid increase in publication volume, along with the diverse contributions from countries, publishers, and organizations, not only highlights current research trends but also lays a robust foundation for future collaborative efforts and advancements in the field.
6 Real-World Applications and Industry Trends
This section is dedicated to addressing RQ-4 by reviewing real-world applications through case studies and discussing approximately ten AI and robotics companies in agriculture.
6.1 Case Studies